OpenStudy (anonymous):

Factor the polyomial completely using any method 12x^3-6x^2+2x-1

OpenStudy (anonymous):

\[6x^2(2x -1) +2x-1= 0 \] \[(2x-1)(6x^2 +1)=0\]

jhonyy9 (jhonyy9):

6x2(2x-1)+2x-1=(6x2+1)(2x-1)

OpenStudy (anonymous):

12x^3-6x^2+2x-1 6x^2(2x-1)+1(2x-1) (6x^2+1)(2x-1)

OpenStudy (anonymous):

Thanks